Southside Pizzeria wants to improve its ability to manage the ingredient costs associated with making and selling its pizzas. For the month of June, the company plans to make 1,000 pizzas. It has created a planning budget that includes a cost formula for mozzarella cheese of $2.40 per pizza. At the end of June, Southside actually sold 1,100 pizzas and the actual cost of the cheese that it used during the month was $2,632. Required 1. What is the mozzarella cheese activity variance for June? 2. What is the mozzarella cheese spending variance for June? 3. Assume that the company establishes a price standard of $0.30 per ounce for mozzarella cheese and a quantity standard of eight ounces of cheese per pizza. Also, assume that Southside actually used 9,400 ounces of cheese during the month to make 1,100 pizzas. a. What is the materials price variance for mozzarella cheese for June? b. What is the materials quantity variance for mozzarella cheese for June? . What is the materials spending variance for mozzarella cheese for June? Complete this question by entering your answers in the tabs below. Required 1Required 2 Required 3 What is the mozzarella cheese activity variance for June? (Indicate the effect of each variance by selecting F for favorable, U for unfavorable, and None for no effect (i.e., zero variance).) ctivity Variance Required 1 Required 2Complete this question by entering your answers in the tabs below. Required 1Required 2 Required 3 What is the mozzarella cheese spending variance for June? (Indicate the effect of each variance by selecting F for favorable U for unfavorable, and None for no effect (i.e., zero variance).) Spending Variance Required 1 Required 3Complete this question by entering your answers in the tabs below. Required 1 Required 2Required 3 Assume that the company establishes a price standard of $0.30 per ounce for mozzarella cheese and a quantity standard of eight ounces of cheese per pizza. Also, assume that Southside actually used 9,400 ounces of cheese during the month to make 1,100 pizzas. a. What is the materials price variance for mozzarella cheese for June? b. What is the materials quantity variance for mozzarella cheese for June? c. What is the materials spending variance for mozzarella cheese for June? (indicate the effect of each variance by selecting F for favorable, U for unfavorable, and None for no effect (i.e., zero variance).) Show less a. Materials price variance b. Materials quantity variance c. | Materials spending variance Required 2 Required 3
Answer #1

1) Activity variance = (1000-1100)*2.4 = $240 U

2) Spending variance = (1100*2.4)-2632 = 8 F

3) Material price variance = (0.30*9400)-2632 = $188 F

Material quantity variance = (1100*8-9400)*0.30 = 180 U

Total material variance = (1100*.30*8)-2632 = $8 F